OpenStudy (anonymous):

Solve the differential equation: (dy/dx)=(2y/x)

OpenStudy (anonymous):

\[\frac{1}{2y}dy = \frac{1}{x}dx \] Separable differential equation.

OpenStudy (anonymous):

and then the antiderivative of (1/2y)dy = the antiderivative of (1/x)dx

OpenStudy (anonymous):

yes, you will get something logarithmic on both sides, usually they want you to term it as a function of y again.

OpenStudy (anonymous):

okay, i got to: (1/2)lny= lnx+ C

OpenStudy (anonymous):

\[ \frac{1}{2}\ln y = \ln x +C = \ln \sqrt{y} \]

OpenStudy (anonymous):

Not sure if they want you to do that though, but you could use exponential properties now to get rid of the natural log terms and then square the entire solution.

OpenStudy (anonymous):

so that leaves me with e(lnx+c) = e(ln sqrt y) x+ Ce = sqrt y

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[\Large e^{\ln x} = x\]

OpenStudy (anonymous):

\[\large \sqrt{y}=x+e^c \]

OpenStudy (anonymous):

Note the e^C is usually just denoted as another constant, so you can write C again.

OpenStudy (anonymous):

\[ \large y=x^2+2e^{c}x+e^{2c} \] I would keep it like that, or you can substitute arbitrary constants, usually they accept it like that.

OpenStudy (anonymous):

ah okay. so then I get y=x^2 + C

OpenStudy (anonymous):

i'm not familiar with how wolfram alpha denotes C http://www.wolframalpha.com/input/?i=solve+differential+equation+%28dy%2Fdx%29%3D%282y%2Fx%29

OpenStudy (anonymous):

let me check.

OpenStudy (anonymous):

is x^2 + c equilavent to their answer?

OpenStudy (anonymous):

yes I see that, well I guess wolfram ignores the constants we have derived above, and just takes y=cx^2 as the most general solution, it's not wrong when you write it as we did it above though.

OpenStudy (anonymous):

yes it is.

OpenStudy (anonymous):

okay that helps, thanks a lot

OpenStudy (anonymous):

you're welcome.